### v2

@@ Abstract
Despite the wide availability of COVID-19 vaccines in the United States and their effectiveness in reducing hospitalizations and mortality during the pandemic, a majority of Americans chose not to be vaccinated during 2021. Recent work shows that vaccine misinformation affects intentions in controlled settings, but does not link it to real-world vaccination rates. Here, we present observational evidence of a causal relationship between exposure to antivaccine content and vaccination rates, and estimate the size of this effect. We present a compartmental epidemic model that includes vaccination, vaccine hesitancy, and exposure to antivaccine content. We fit the model to data to determine that a geographical pattern of exposure to online antivaccine content across US counties explains reduced vaccine uptake in the same counties. We find observational evidence that exposure to antivaccine content on Twitter caused about 14,000 people to refuse vaccination between February and August 2021 in the US, resulting in at least 510 additional cases and 8 additional deaths. This work provides a methodology for linking online speech with offline epidemic outcomes. Our findings should inform social media moderation policy as well as public health interventions.
@@ 3 Antivaccine Tweets Prevent Vaccinations
We can use measurements of vaccine effectiveness to provide a lower bound for the number of COVID cases and deaths that may have resulted from this reduction in vaccination. We estimate that among the 14,086 people who remained unvaccinated as a result of antivaccine tweets, there were about 511 COVID cases and 8 COVID-attributable deaths during February-August 2021 that would have been prevented without this additional vaccine hesitancy (see Methods). This represents a lower bound on the total impact because there would have been secondary infections outside the vaccine-hesitant population. Additionally, there would have been more cases and deaths following August 2021, which are not counted here.
@@ 5.5 Effect Estimations
We calculate the number of people across the whole US who remained unvaccinated as a result of exposure to antivaccine tweets to be 14,086 (see Supplementary Methods). We then wish to estimate how many cases and deaths may have been prevented among these people if they had been vaccinated instead. Assuming their infection and death rates were typical for unvaccinated people during the February-August 2021 time period (3,870 cases and 57 deaths per 100,000 people, see Supplementary Methods), we estimate that about 511 cases and 8 deaths occurred in this population. Of these cases and deaths, approximately 93% of cases and 94% of deaths may have been prevented by vaccination [31, 32], yielding the numbers of COVID cases and deaths attributable to antivaccine tweets, as reported in our results.
@@ Acknowledgements
We are grateful to Marissa Donofrio for annotating tweets, to Bjarke Mønsted and Sune Lehmann for sharing additional annotation data, and to Vincent Jansen, Alessandro Flammini, and Yong-Yeol Ahn for helpful discussion. We also thank two anonymous reviewers who provided helpful feedback, highlighting an error in an earlier version of the manuscript.
@@ A.5 Deriving the ATE Estimand
We also computed this quantity in an alternative way by simulating the SIRVA model with the inferred parameters in a counterfactual scenario where the exposure in all counties was set to zero. Comparing the real data with this scenario allowed us to infer that exposure accounted for 980,000 fewer vaccinations. This alternative estimate is compatible with that obtained with the ATE-based method and helps validate our assumption about the small effect size. However, the ATE-based method is better able to explicitly account for potential confounding factors.
@@ A.6 Estimating Case and Death Rates among the Unvaccinated Population
Computing these risks for an unvaccinated person in the United States over the period February-August 2021 using CDC data for cases, deaths, and vaccinations and values of $\lambda_{C}=0.93$ and $\lambda_{D}=0.94$ [31, 32], we find $P(C\mid\bar{V})=0.0387$ and $P(D\mid\bar{V})=0.00057$. This equates to 511 cases and 8 deaths per 100,000 unvaccinated people, which we use in the main text to estimate the numbers of cases and deaths resulting from exposure to antivaccine content on Twitter.

### v3

@@ Abstract
Despite the wide availability of COVID-19 vaccines in the United States and their effectiveness in reducing hospitalizations and mortality during the pandemic, a majority of Americans chose not to be vaccinated during 2021. Recent work shows that vaccine misinformation affects intentions in controlled settings, but does not link it to real-world vaccination rates. Here, we present observational evidence of a causal relationship between exposure to antivaccine content and vaccination rates, and estimate the size of this effect. We present a compartmental epidemic model that includes vaccination, vaccine hesitancy, and exposure to antivaccine content. We fit the model to data to determine that a geographical pattern of exposure to online antivaccine content across US counties explains reduced vaccine uptake in the same counties. We find observational evidence that exposure to antivaccine content on Twitter caused about 14,000 people to refuse vaccination between February and August 2021 in the US, resulting in at least 545 additional cases and 8 additional deaths. This work provides a methodology for linking online speech with offline epidemic outcomes. Our findings should inform social media moderation policy as well as public health interventions.
@@ 3 Antivaccine Tweets Prevent Vaccinations
We can use measurements of vaccine effectiveness to provide a lower bound for the number of COVID cases and deaths that may have resulted from this reduction in vaccination. We estimate that among the 14,086 people who remained unvaccinated as a result of antivaccine tweets, there were about 545 COVID cases and 8 COVID-attributable deaths during February-August 2021 that would have been prevented without this additional vaccine hesitancy (see Methods). This represents a lower bound on the total impact because there would have been secondary infections outside the vaccine-hesitant population. Additionally, there would have been more cases and deaths following August 2021, which are not counted here.
@@ 5.5 Effect Estimations
We calculate the number of people across the whole US who remained unvaccinated as a result of exposure to antivaccine tweets to be 14,086 (see Supplementary Methods). We then wish to estimate how many cases and deaths may have been prevented among these people if they had been vaccinated instead. Assuming their infection and death rates were typical for unvaccinated people during the February-August 2021 time period (3,870 cases and 57 deaths per 100,000 people, see Supplementary Methods), we estimate that about 545 COVID cases and 8 deaths that occurred in this population are attributable to antivaccine tweets, as reported in our results.
@@ Acknowledgements
We are grateful to Marissa Donofrio for annotating tweets, to Bjarke Mønsted and Sune Lehmann for sharing additional annotation data, and to Vincent Jansen, Alessandro Flammini, and Yong-Yeol Ahn for helpful discussion. We also thank four anonymous reviewers who provided helpful feedback, highlighting an error in an earlier version of the manuscript.
@@ A.5 Deriving the ATE Estimand
We also computed this quantity in an alternative way by simulating the SIRVA model with the inferred parameters in a counterfactual scenario where the exposure in all counties was set to zero. Comparing the real data with this scenario allowed us to infer an alternative estimate compatible with that obtained with the ATE-based method, helping validate our assumption about the small effect size. However, the ATE-based method is better able to explicitly account for potential confounding factors.
@@ A.6 Estimating Case and Death Rates among the Unvaccinated Population
Computing these risks for an unvaccinated person in the United States over the period February-August 2021 using CDC data for cases, deaths, and vaccinations and values of $\lambda_{C}=0.93$ and $\lambda_{D}=0.94$ [43, 44], we find $P(C\mid\bar{V})=0.0387$ and $P(D\mid\bar{V})=0.00057$. This equates to 3,870 cases and 57 deaths per 100,000 unvaccinated people, which we use in the main text to estimate the numbers of cases and deaths resulting from exposure to antivaccine content on Twitter.